OpenStudy (anonymous):

Prove that there exists a pair of consecutive integers such that one of these integers is a perfect square and the other is a perfect cube.

ganeshie8 (ganeshie8):

there is a trivial example : 0 = 0^3 1 = 1^2

Parth (parthkohli):

8, 9 too.